Average collection cycle. Use the following account information for Rian Company.

2013 and 2014 Selected Balance Sheet Accounts of Rian Company

12/31/14 . 12/31/13 Change

Accounts Recievable $33, 299 .    $40,466    $7,147

Inventory $49,815 $52,841 $3,026

Accounts Payable $24,930 $23,111 $1,819

2014 Selected Income Statement Items for Rian Company

Cash sales

​$416,000

Credit sales

​$460,000

Total sales

​$876,000

Cost of goods sold

​$501,293

Rian Company had set a target of 27.4 days for the collection cycle for 2014. If total sales had remained at $876,000​ how much of the sales revenue would have needed to be cash sales for the company to have met the collection​ goal?

If total sales had remained at ​$876,000​, how much of the sales revenue would have needed to be cash sales for the company to meet the collection​ goal?
Answer #1

Average Receivable Turnover = Net Credit Sales / Average Accounts Receivable

Average Accounts Receivable = (33299+40466+7147)/3 = 26970.67

Collection Cycle = 27.4 Days = 365 / Average Receivable Turnover

=> Average Receivable Turnover = 365/27.4=13.32117

Average Receivable Turnover = 13.32117 Days = Credit Sales / 26970.67

=>Credit Sales = 26970.67*13.32117 = 359280.9

Cash Sales = Total Sales – Credit Sales = $876,000 - 359280.9= $ 516719.1.